a rural population (given in thousands) is thought to decline according to the equation p=15e^(-0.1t). if t=0 at the beginning of 1998. calculate the numbers in the population at the beginning of 1990,2000,2016
1
Expert's answer
2012-08-31T09:33:26-0400
1900 -- 33,38311
1900 – 1998 = -8 year
p=15e^(-0.1*(-8))=33,383

2000---12,28096
2000 – 1998 = 2 year
p=15e^(-0.1*(2))= 12,28096

2016 ----2,479483
2016 – 1998 = 18 year
p=15e^(-0.1*(18))= 2,479483